OpenStudy (anonymous):

There are 35 questions on Allison’s English test. The multiple choice questions are worth 2 points each, and the short answer questions are worth 5 points each. The entire test is worth 100 points. Let multiple question be x. And short answers be y. How many short answer questions are on the test?

OpenStudy (anonymous):

\[\large x+y=35\\\large 2x+5y=100\]

OpenStudy (anonymous):

2x + 5y = 100 || 35 = x + y 35 = x + y 35 - y = x SUB... 2x + 5y = 100 2(35-y) + 5y = 100 70 - 2y + 5y = 100 3y = 30 y = 10 SUB BACK 35 = x + y 35 =x + 10 x = 25

OpenStudy (anonymous):

Thanks guys